Your Uncle who lives in Kintinku Ward in remote Singida region is a typical entrepreneur. He owns a retail shop selling general merchandise products. He however plans to buy a TV set and a satellite dish so that football funs can come to his place to view soccer matches at a fee. He also plans to sell cold drinks to customers in general. He is aware that electrical power will be required though his village is not served by TANESCO. To solve this power problem, he has two options.

Option A requires him to go solar. A 150Wp solar PV system is needed. This system is capable of powering ten direct current fluorescent light bulbs, a radio cassette, a 21-inch colour television and a video player (the system would alternatively also power mobile phone charging facility and hair cutting saloon would suffice. Such a solar PV would require the following components and expenditures:

Item

Amount

First cost (TAS)

Annual cost/revenue

(TAS)

Annual cost/revenue increase per year TAS

Life span - Years

Salvage value (TAS)

Solar Module 75W

2

1,160,000

15

100,000

Charge regulator 20W

1

120000

15

-

Solar batteries

2

200,000

3

20,000

Inverter DC/AC, 500W Trace

1

850,000

3

-

Energy saver bulbs 12VDC

10

150000

75,000

5,000

15

-

Installation & wiring materials, SET

1

150000

15

-

Transport charges

1

200000

Labour charge, technicians for installation

1

200000

Maintenance costs

15,000

1,200

15

-

Expected revenue

7,000,000

500,000

Option B would need purchasing 1 KV diesel generator. Such a generator would require the following components and expenditures (See table below):

Item

Amount

First cost (TAS)

Annual cost/revenue

(TAS)

Annual cost/revenue increase per year TAS

Life span - Years

Salvage value (TAS)

I KVA generator

1

1,000,000

8

100,000

Diesel

720 litres per annum.

1,152,000

200,000

-

-

Maintenance costs

50,000

10,000

Revenue

6,000,000

400,000
Answer #1

Option 1:

Where,

Initial cost = Sum of all the equipments initial cost of 15 year life multiplied by their amouts.

Annual revenue = 75000 + 700000 = 7075000 and increasing by 520000 each year (i.e.5000+15000+500000)

Salvage value = Multiplied by amount

CF each year = Annual revenue + salvage value -Initial cost

Discounting CF to persent worth by assuming a rate of 10%.

PW = \sum \frac{CF_{n}}{(1+R)^{n}}

Years Initial cost Salvage value Annual rev CF Discounted CF
0 25490000 -25490000 -25490000
1 7075000 7075000 6431818.182
2 7595000 7595000 6276859.504
3 8115000 8115000 6096919.609
4 8635000 8635000 5897821.187
5 9155000 9155000 5684534.713
6 9675000 9675000 5461285.273
7 10195000 10195000 5231647.015
8 10715000 10715000 4998626.579
9 11235000 11235000 4764736.742
10 11755000 11755000 4532061.367
11 12275000 12275000 4302312.616
12 12795000 12795000 4076881.313
13 13315000 13315000 3856881.216
14 13835000 13835000 3643187.903
15 200000 14355000 14555000 3484351.279
PW 49249924.5

Now, computing the 3 years of life equipments separately, we get the following PW

Years Initial cost Salvage CF Discounted
0 1250000 -1250000 -1250000
1 0
2 0
3 40000 40000 53240
PW -1196760

Now adding both the PW together, we get the PW of the option as,

Total PW 48053164

Option 2;

Where, initial cost = Sum of all the equipment's initial cost

Revenue = sum of all equipment, increasing by the increasing cost of all the equipment (i.e. 200000+10000+400000)

CF = Revenue + Salvage - initial cost

Discounting by 10% - PW = \sum \frac{CF_{n}}{(1+R)^{n}}

PW = 47246278.87

Year Initial cost Revenue Salvage value CF Discounted CF
0 1000000 -1000000 -1000000
1 7202000 7202000 6547272.727
2 7812000 7812000 6456198.347
3 8422000 8422000 6327573.253
4 9032000 9032000 6168977.529
5 9642000 9642000 5986923.397
6 10252000 10252000 5786986.731
7 10862000 10862000 5573923.48
8 11472000 100000 11572000 5398423.404
PW 47246278.87

Thus Option 1 has higher PW, thus we go with option 1.